OpenStudy (anonymous):

Find the probability. An IRS auditor randomly selects 3 tax returns (without replacement) from 47 returns of which 5 contain errors. What is the probability that she selects none of those containing errors? A. 0.0012 B. 0.0006 C. 0.708 D. 0.7136

OpenStudy (anonymous):

\[\frac{42}{47}\times \frac{41}{46}\times \frac{40}{45}\]whatever that is

OpenStudy (anonymous):

reasoning that if 5 contain errors then 42 do not, then as before pick one without an error, leaves 46 to choose from of which 41 do not contain errors etc

OpenStudy (anonymous):

i know you need the decimal right? http://www.wolframalpha.com/input/?i=42%2F47*41%2F46*40%2F45